Consider the polynomial of a complex variable z given by f(z)=z³-8z²+25z-26



(a) Show that z-3+2i is factor of f(z)



(b) Develop the quadratic factor of f(z)



(c) Hence, completely solve the equation f(z)=0

1
Expert's answer
2021-11-08T06:10:44-0500

Consider the polynomial of a complex variable "z" given by "f(z)=z^3-8z^2+25z-26."


(a) Let us show that "z-3+2i" is factor of "f(z)." Since

"f(3-2i)=(3-2i)^3-8(3-2i)^2+25(3-2i)-26\\
=27-54i+36i^2-8i^3-8(9-12i+4i^2)+75-50i-26\\
=27-54i-36+8i-72+96i+32+75-50i-26=0,"

we conclude that "3-2i" is a root of "f," and hence according to Bezout Theorem, "z-3+2i" is factor of "f(z)."


(b) Let us develop the quadratic factor of "f(z)." Since the coefficients of "f" are the real numbers, "3+2i" is also a root of "f," and hence "z\u22123-2i" is factor of "f(z)." It follows that "(z\u22123+2i)(z\u22123-2i ) =(z-3)^2-4i^2=z^2-6z+9+4=z^2-6z+13"

is the quadratic factor of "f(z)."



(c) Hence, let us completely solve the equation "f(z)=0."

It follows that

"f(z)=z^3-8z^2+25z-26=(z^2-6z+13)(z-2),"

and thus "z=2" is the third root of the equation.

We conclude that the roots of the equations are:

"3-2i,\ 3+2i" and 2.